Calculate the pH of 0.12 M NH3 solution. Kb NH3 = 1.8 x10-5 


1
Expert's answer
2021-08-18T13:18:04-0400

Solution:

Ammonia (NH3) is a weak base.

The equilibrium equation for its reaction with water:

NH3(aq) + H2O(l) ⇌ NH4+(aq) + OH(aq)


From this, we can construct an ICE (Initial, Change, Equilibrium) Table:




The Kb expression for NH3 is:



Kb = [ x ] × [ x ] / [ 0.12 - x ] = 1.8×10-5


C(NH3) / Kb = (0.12) / (1.8×10-5) = 6666.67 >> 500

Since the ratio is greater than 500, we can remove (–x) from the equation.


The equation simplifies to:

Kb = [ x ] × [ x ] / [ 0.12 ]

Kb = x2 / 0.12

Solving for x:

x2 = (1.8×10-5) × 0.12 = 2.16×10-6

x = (2.16×10-6)0.5 = 0.00147


This means [OH] = [NH4+] = x = 0.00147 M


From [OH], we can calculate pOH:

pOH = -log[OH-] = -log(0.00147) = 2.83

 

From pOH, we can calculate pH:

pH + pOH = 14;

pH = 14 - pOH = 14 - 2.83 = 11.17

pH = 11.17

 

Answer: pH = 11.17